OpenStudy (anonymous):

Solve for a in the inequality below. –30 + 5a < =16

OpenStudy (gorv):

add 30 on both side

OpenStudy (gorv):

and after that divide both side by 5

OpenStudy (gorv):

@eddiet14

OpenStudy (anonymous):

Simplifying -30 + 5a = 16 Solving -30 + 5a = 16 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '30' to each side of the equation. -30 + 30 + 5a = 16 + 30 Combine like terms: -30 + 30 = 0 0 + 5a = 16 + 30 5a = 16 + 30 Combine like terms: 16 + 30 = 46 5a = 46 Divide each side by '5'. a = 9.2 Simplifying a = 9.2

OpenStudy (anonymous):

a=46/5 you have to divide that

OpenStudy (anonymous):

got it thank you guys!

OpenStudy (gorv):

welcome:)